Can you please solve this problem and show me how to do it? Than you.
How many g of gold should a coin of 53% gold be if when combined with a 5 g pure gold necklace, it forms a metal that is 76% gold?

Answer by stanbon(75887)

How many g of gold should a coin of 53% gold be if when combined with a 5 g pure gold necklace, it forms a metal that is 76% gold?
------
Equation:
gold + gold = gold
0.53x + 1*5 = 0.76(x+5)
------
Multiply thru by 100 to get::
53x + 100*5 = 76*x + 76*5
------
23x = 24*5
x = 5.217 grams (amt of 53% gold to add)
---------------
Cheers,
Stan H.